# Phenome-Wide Multi-Omics Integration Uncovers Distinct Archetypes of Human Aging

### v1

@@ 1 Introduction
Here we leveraged a large-scale, well-characterized human cohort (Human-Phenotype Project, HPP) [25] consisting of 10K longitudinal study, which includes comprehensive clinical, physiological, behavioral, environmental, and multi-omics profiling, to investigate the complex dynamics of human aging. By integrating high-dimensional data spanning the transcriptome, lipidome, metabolome, and microbiome, we developed and rigorously validated a novel multi-omics aging clock. Most existing aging clocks (like epigenetic clocks) produce one output, however, our work goes beyond providing a single aging score and provide holistic and biologically informative picture of systemic decline such as risk association. Through unsupervised clustering of the integrated omics landscape, we identified distinct molecular subtypes, thereby uncovering the profound molecular and phenotypic heterogeneity inherent to the human aging process. Subsequent pathway enrichment and network analyses of these subtypes enabled us to deconstruct the biological consequences of aging and demonstrate how key biological processes are dynamically reweighted across divergent aging trajectories.
@@ 2.2 Multi-omics aging clocks predict biological age
We developed a multi-omics aging clock using five types of omics data derived from human PBMC, feces, and saliva samples. During the training phase, we compared three machine learning methods (least absolute shrinkage selection operator (LASSO) [26], Elastic Net [27], LightGBM [28]) to train multi-omics age clock models. We evaluated the performance of the aging clock by comparing the predicted age with the CA (Pearson’s R, R2, root mean square error (r.m.s.e.), median absolute error (m.a.e.) and median absolute error (med.a.e)). A cross-validation step was implemented, where 80% of the samples were randomly selected for model training, while the remaining 20% of the data were used as an internal validation set. In our initial experiments, we found that LightGBM showed superior performance across all omics types. Consequently, we selected it as the final model and used SHapley Additive exPlanations (SHAP) [29] to identify omics features associated with predicting chronological age. The strong correlation between chronological age and predicted transcriptional age demonstrated the accuracy of our aging clock (Fig. 2a). We further confirmed the robustness of the aging clock model through 10-fold cross-validation, which showed only slight performance changes in all metrics.
@@
To evaluate the physiological relevance of age acceleration across biological strata, we assessed associations between five omics-based age-acceleration (AgeAccel) scores and a comprehensive panel of parameters spanning 12 physiological and lifestyle domains (Fig. 2). AgeAccel scores, computed as residuals from regressing biological age on chronological age, differed substantially in their links to systemic health, after adjusting for multiple hypotheses (false discovery rate (FDR) << 0.05). Of these, lifestyle was significantly associated with all five omics-based AgeAccel metrics (Proportion=0.10 for transcriptomics, Proportion=0.01 for metabolomics, Proportion=0.28 for lipidomics, Proportion=0.78 for gut microbiome and Proportion=0.77 for oral microbiome). Gut-microbiome AgeAccel showed the broadest association profile (significant associations with Liver health, Insulin resistance, frailty, lifestyle and hematopoietic system) and was the only omics-derived AgeAccel associated with frailty,onsistent with previous findings by [30]. Oral-microbiome AgeAccel showed a similar but weaker pattern. The oral-microbiome AgeAccel showed the strongest association with DXA-derived bone density (Proportion=0.24), independent of sex and BMI. This pattern aligns with previous reported links between periodontal dysbiosis and systemic bone loss/osteoporosis [31], as well as with elevated pro-inflammatory cytokines promoting osteoclastogenesis [32, 33]. Moreover, oral microbiome composition significantly shifts (p = XX) with age and reflects aspects of biological aging, providing a plausible rationale for an AgeAccel–bone density association, which is aligned with [34]. In contrast to the broad profiles of the microbiome-based metrics, the remaining molecular AgeAccel clocks exhibited more circumscribed physiological associations. The Transcriptomic-AgeAccel was predominantly linked to metabolic dysregulation, with biomarkers of insulin resistance constituting nearly half of its significant associations (Proportion=0.45). Weaker, yet significant, links were also observed with diet (Proportion=0.11), sleep (Proportion=0.11), and lifestyle (Proportion=0.10), while its connection to the cardiovascular system was minimal (Proportion=0.02). The Metabolomic-AgeAccel displayed an even more specific signature, being almost exclusively associated with the renal function system (Proportion=0.50). Conversely, the Lipidomic-AgeAccel captured a more pleiotropic signal; it was most strongly dominated by associations with diet (Proportion=0.67) but was also substantially linked to the hematopoietic (Proportion=0.43) and immune (Proportion=0.40) systems, in addition to lifestyle (Proportion=0.28).
@@
In summary, our multi-omics aging clock, constructed using a nonlinear model, accurately predicted biological age, suggesting that aging-related molecular and microbial changes follow nonlinear patterns. Furthermore, using multi-omics data, we quantify age acceleration at different biological layers. While molecular and microbial profiles exhibited some consistent association with phenotypic changes during aging acceleration, we observed substantial divergent associations across multiple systems-level categories, such as bone density, insulin resistance, frailty, and the hematopoietic system, highlighting the significance of integrating multi-omics data for constructing a comprehensive aging clock.
@@ 2.3 Systemic Signatures and Molecular Dynamics of Aging Phenotypes
Figure 3: Distinct molecular aging trajectories and systemic signatures in accelerated and decelerated aging. a, c, Sankey diagrams illustrating the dynamics of significantly altered biological features in the accelerated (a) and decelerated (c) aging groups. The diagrams show the number of features that significantly increase or decrease in subsequent age brackets relative to the 40–45 age group baseline. Significance was assessed using a two-sided Mann-Whitney U-test (FDR-adjusted p << 0.05). b, d Pathway enrichment analysis of genes significantly altered between ages 55–60 and 60–65 in the accelerated (b) and decelerated (d) aging groups. e, f, Temporal trajectory clusters of all multi-omics features for the accelerated (e) and decelerated (f) aging groups, identified using the Fuzzy C-means algorithm. Each panel represents a distinct cluster, showing the z-scored trajectories of individual molecules (thin lines) and the cluster centroid (thick line) across age. The number of features (n) in each cluster is indicated. g, Heatmap showing the percentage of features that are significantly different between the accelerated and decelerated aging groups across various physiological systems. Color intensity corresponds to the percentage of significant features within each system, identified by one-way ANOVA.
@@
To understand the biological processes that distinguish accelerated from decelerated aging, we first classified participates into aging-accelerated and aging-decelerated groups, and further categorized them into distinct age stages according to their ages and investigated the altered biological features within each age stage compared to the baseline (40–45 years old; Methods). Our analysis revealed that accelerated aging is marked by an early and large-scale disruption of molecular and microbial homeostasis. A substantial number of features showed significant changes from the 40–45 age baseline as early as the 50–55 age bracket, with these alterations accumulating progressively into later life (Fig. 3a). In stark contrast, the decelerated aging group exhibited a blunted and delayed molecular aging phenotype, with a significantly smaller number of features changing in early mid-life while the number of features that change during the transition from middle age to old age increases dramatically. (Fig. 3c). Next, we performed Gene Ontology (GO) [39] enrichment analysis on transcriptomic data from individuals aged 55–60 versus 60–65, identifying pathways in which genes that were individually non-significant collectively reached significance (Methods). We identified some functional modules that were reported in previous studies, but they played different roles in different aging patterns during human aging. For instance, in accelerated aging, enrichment for T-cell activation [40, 41] indicates an adaptive-immune, adhesion-driven inflammatory state that is consistent with chronic antigenic stimulation and tissue infiltration (Fig. 3b). By contrast, decelerated aging is marked by enrichment of regulation of small-molecule metabolic process, aligning with preserved metabolic and redox homeostasis that supports genome maintenance and antigen presentation while restraining excessive immune activation[42, 43] (Fig. 3d).
@@ 4.2 Body system-derived features
To systematically analyze the interrelationships among physiological modules, we organized the hundreds of measurement parameters and questionnaire information collected at baseline into the following core system-level categories. The specific characteristics, available sample size, and detailed description of each category are shown in Extended Data Table 1.
@@
4.3 Transcriptomics preprocessing
Transcriptomic analysis was performed on mononuclear cells (PBMCs) isolated from peripheral blood. Total RNA was extracted using the QIAGEN RNeasy Mini Kit and the manufacturer’s instructions were strictly followed. Transcriptome sequencing libraries were constructed using the Illumina TruSeq Stranded Total RNA Library Prep Kit with Ribo-Zero Globin. In brief, the process first efficiently removes ribosomal RNA (rRNA) and globin messenger RNA (globin mRNA) from total RNA, and then the purified RNA is fragmented. Reverse transcription is performed using random primers with Illumina-specific adapters to synthesize the first-strand cDNA. The second strand is then synthesized and ligated to double-end adapters, and the final cDNA library is amplified by PCR to achieve sufficient concentration for loading. The prepared library was evaluated for fragment size distribution using an Agilent Bioanalyzer and accurately quantified using the Qubit fluorescence quantification method (Thermo Fisher Scientific). After equimolar pooling, the validated libraries were sequenced on the Illumina NovaSeq 6000 platform in a paired-end 150 bp (2x150 bp) sequencing mode. On average, each sample generated about 50 million sequencing reads. Image analysis and base calling of raw sequencing data were performed by Illumina standard processes. We first used FastQC to assess the quality of raw reads, and then used Trimmomatic to remove adapter sequences and low-quality bases. The clean reads that passed quality control were aligned to the GRCh38 (hg38) human reference genome using STAR (Spliced Transcripts Alignment to a Reference) alignment software. Subsequently, the featureCounts tool was used to count the reads in the exon region of each gene to quantify the gene expression level. In the data preprocessing stage, we filtered the gene expression matrix. In all samples, low-expression genes with counts per million (CPM << 1) in at least half of the samples were removed. The filtered data were then normalized using the TMM (Trimmed Mean of M-values) method in the edgeR software package to correct for differences in library size between samples for downstream differential expression and association analysis.
4.4 Metabolomics preprocessing
Metabolic profiling was performed on serum samples using the high-throughput nuclear magnetic resonance (NMR) spectroscopy platform from Nightingale Health, Finland. This platform provides quantitative measurements for over 200 metabolic biomarkers from a single sample. Sample preparation was automated using robotic liquid handlers, where a standardized volume of each sample was mixed with a buffered saline solution and transferred into 96-well plates. Data acquisition was performed on multiple NMR spectrometers operating under strictly controlled conditions, including standardized temperature. For each sample, a one-dimensional (1D) proton (¹H) NMR spectrum was acquired. The resulting raw spectral data were processed and quantified using Nightingale’s proprietary computational pipeline. This process involved several steps, including phase and baseline correction, followed by the deconvolution of the NMR spectrum to calculate the signal area for each metabolic biomarker. The absolute concentration (e.g., in mmol/L) of each biomarker was then determined by comparing its signal area to that of specific reference standards with known concentrations. A rigorous, multi-level quality control (QC) process was embedded throughout the workflow. Certified control samples were analyzed alongside the study samples to monitor and validate the quantification process. Key instrument performance metrics, such as signal-to-noise ratio, line width, and water suppression efficiency, were continuously monitored. Any sample that failed to meet the predefined QC criteria was flagged and subject to re-analysis, ensuring high data quality and consistency across the cohort.
4.5 Serum lipidomics preprocessing
Comprehensive lipidomic profiling was performed on serum samples using an ultra-performance liquid chromatography-high resolution mass spectrometry (UPLC-HRMS) platform. Lipids were extracted from serum using a liquid-liquid extraction method, typically with a chloroform:methanol based solvent system. A cocktail of non-endogenous lipid species, representing various lipid classes, was added as internal standards prior to the extraction to control for analytical variability and enable accurate quantification. Following extraction, the lipid-containing organic phase was evaporated to dryness under a nitrogen stream and reconstituted in an appropriate solvent mixture for injection. The lipid extract was then injected into a UPLC system equipped with a reversed-phase column (e.g., C18) for chromatographic separation. The mobile phases consisted of an aqueous solution and an organic mixture (e.g., acetonitrile/isopropanol), both containing additives such as ammonium formate to enhance ionization. The column eluent was introduced into a high-resolution mass spectrometer (e.g., a Thermo Q Exactive series or Agilent Q-TOF instrument). Data were acquired in both positive and negative ionization modes to ensure broad coverage of diverse lipid classes. Full scan mass spectra were acquired, and data-dependent MS/MS fragmentation data were subsequently collected for the most abundant precursor ions to facilitate structural elucidation. The raw data files were processed using a specialized software pipeline (e.g., MS-DIAL or LipidSearch). The workflow included peak picking, retention time alignment, and deconvolution. Lipid species were identified based on two key criteria: accurate mass matching of the precursor ion (typically with a mass error ¡5 ppm) and comparison of the experimental MS/MS fragmentation pattern against established spectral libraries and public databases (e.g., LIPID MAPS). The peak area of each identified lipid was normalized to the peak area of its corresponding class-specific internal standard for relative or absolute quantification. The final dataset was filtered based on quality control criteria, including high detection frequency and low coefficient of variation (CV) in pooled QC samples, to ensure data robustness.
4.6 Microbiome preprocessing
Microbiome composition was profiled from fecal and saliva samples using 16S rRNA gene amplicon sequencing. DNA was extracted from fecal samples and from saliva collected in standardized devices (e.g., Oragene DNA kits) using commercial kits optimized for each respective sample type (e.g., DNeasy PowerSoil Kit, QIAGEN). For amplification, the V3-V4 and V4 hypervariable regions of the 16S rRNA gene were targeted for fecal and oral samples, respectively, using region-specific primers flanked by Illumina sequencing adapters. The purified amplicons were then pooled and sequenced on an Illumina platform (e.g., MiSeq or NovaSeq), generating paired-end reads. The raw sequencing data for both microbiome datasets underwent a unified and rigorous bioinformatics pipeline using QIIME 2 or the DADA2 package. This process included quality filtering, trimming of low-quality bases and primers, denoising to correct sequencing errors, merging of paired-end reads, and chimera removal. This workflow generated a high-resolution feature table of Amplicon Sequence Variants (ASVs) for each dataset. Taxonomic classification of ASVs was performed by aligning representative sequences against curated reference databases chosen for their specificity to the sample source. Specifically, the Greengenes or SILVA database was used for the gut microbiome data, while the specialized Human Oral Microbiome Database (HOMD) was used for the oral microbiome data to ensure accurate species-level assignment. The final outputs, consisting of ASV abundance tables and their corresponding taxonomic assignments, served as the basis for downstream ecological and statistical analyses, including the calculation of alpha and beta diversity metrics.

### v2

@@ 1 Introduction
Here we leveraged a large-scale, well-characterized human cohort (Human-Phenotype Project, HPP) [25] consisting of 10K longitudinal study, which includes comprehensive clinical, physiological, behavioral, environmental, and multi-omics profiling, to investigate the complex dynamics of human aging. Most existing aging clocks (like epigenetic clocks) produce one output. By integrating high-dimensional data spanning the transcriptome, lipidome, metabolome, and microbiome, we developed and rigorously validated a novel multi-omics aging clock and provided holistic and biologically informative picture of systemic decline such as risk association. Through unsupervised clustering of the integrated omics landscape, we identified distinct molecular subtypes, thereby uncovering the profound molecular and phenotypic heterogeneity inherent to the human aging process. Subsequent pathway enrichment and network analyses of these subtypes enabled us to deconstruct the biological consequences of aging and demonstrate how key biological processes are dynamically reweighted across divergent aging trajectories.
@@ 2.2 Multi-omics aging clocks predict biological age
We developed a multi-omics aging clock using five types of omics data derived from human PBMC, feces, and saliva samples. During the training phase, we compared three machine learning methods (least absolute shrinkage selection operator (LASSO) [26], Elastic Net [27], LightGBM [28]) to train multi-omics age clock models. We evaluated the performance of the aging clock by comparing the predicted age with the CA (Pearson’s R, R2, root mean square error (r.m.s.e.), median absolute error (m.a.e.) and median absolute error (med.a.e)). A cross-validation step was implemented, where 80% of the samples were randomly selected for model training, while the remaining 20% of the data were used as an internal validation set. Fig. 2a underscored the accuracy of our aging clocks. In our initial experiments, we found that LightGBM showed superior performance across all omics types. Consequently, we selected it as the final model and used SHapley Additive exPlanations (SHAP) [29] to identify omics features associated with predicting chronological age. The strong correlation between chronological age and predicted transcriptional age demonstrated the accuracy of our aging clock. We further confirmed the robustness of the aging clock model through 10-fold cross-validation, which showed only slight performance changes in all metrics (Fig. 2b).
@@
To evaluate the physiological relevance of age acceleration across biological strata, we assessed associations between five omics-based age-acceleration (AgeAccel) scores and a comprehensive panel of parameters spanning 12 physiological and lifestyle domains (Fig. 2c). AgeAccel scores, computed as residuals from regressing biological age on chronological age, differed substantially in their links to systemic health, after adjusting for multiple hypotheses (false discovery rate (FDR) << 0.05). Of these, lifestyle was significantly associated with all five omics-based AgeAccel metrics (propprop=0.10 for transcriptomics, propprop=0.01 for metabolomics, propprop=0.28 for lipidomics, propprop=0.78 for gut microbiome and propprop=0.77 for oral microbiome). Gut-microbiome AgeAccel showed the broadest association profile (significant associations with Liver health, Insulin resistance, frailty, lifestyle and hematopoietic system) and was the only omics-derived AgeAccel associated with frailty,onsistent with previous findings by [30]. Oral-microbiome AgeAccel showed a similar but weaker pattern. The oral-microbiome AgeAccel showed the strongest association with DXA-derived bone density (propprop=0.24), independent of sex and BMI. This pattern aligns with previous reported links between periodontal dysbiosis and systemic bone loss [31, 32, 33]. Moreover, oral microbiome composition significantly shifts (P << 2.2e-16) with age and reflects aspects of biological aging, providing a plausible rationale for an AgeAccel–bone density association, which is aligned with [34]. In contrast to the broad profiles of the microbiome-based metrics, the remaining molecular AgeAccel clocks exhibited more circumscribed physiological associations. The Transcriptomic-AgeAccel was predominantly linked to metabolic dysregulation, with biomarkers of insulin resistance constituting nearly half of its significant associations (propprop=0.45). Weaker, yet significant, links were also observed with diet (propprop=0.11), sleep (propprop=0.11), and lifestyle (propprop=0.10), while its connection to the cardiovascular system was minimal (propprop=0.02). The Metabolomic-AgeAccel displayed an even more specific signature, being almost exclusively associated with the renal function system (propprop=0.50). Conversely, the Lipidomic-AgeAccel captured a more pleiotropic signal; it was most strongly dominated by associations with diet (propprop=0.67) but was also substantially linked to the hematopoietic (propprop=0.43) and immune (propprop=0.40) systems, in addition to lifestyle (propprop=0.28).
@@ 2.3 Systemic Signatures and Molecular Dynamics of Aging Phenotypes
Figure 3: Distinct molecular aging trajectories and systemic signatures in accelerated and decelerated aging. a-d, Dynamics of significantly altered biological features across age in the accelerated (a, b) and decelerated (c, d) aging groups. a, c, Sankey diagrams showing the number of features that significantly increase or decrease with age relative to the 40–45-year baseline (two-sided Mann–Whitney U-test, FDR-adjusted P<<0.05). b, d, Pathway enrichment of genes significantly altered between the 55–60 and 60–65-year groups. e, f, Temporal trajectory clusters of all multi-omics features for the accelerated (e) and decelerated (f) aging groups, identified using Fuzzy C-means clustering. Thin lines represent individual z-scored trajectories, and thick lines denote cluster centroids. The number of features (n) per cluster is indicated. g, Heatmap showing the percentage of features that are significantly different between the accelerated and decelerated aging groups across various physiological systems. Color intensity corresponds to the percentage of significant features within each system, identified by one-way ANOVA.
@@
To understand the biological processes that distinguish accelerated from decelerated aging, we first classified participants into aging-accelerated and aging-decelerated groups based on AgeAccel scores, and further categorized them into distinct age stages according to their ages and investigated the altered biological features within each age stage compared to the baseline (40–45 years old; Methods). Our analysis revealed that accelerated aging is marked by an early and large-scale disruption of molecular and microbial homeostasis. A substantial number of features showed significant changes from the 40–45 age baseline as early as the 50–55 age bracket, with these alterations accumulating progressively into later life (Fig. 3a). We then focused on features that were not significantly altered at 55–60 years but became significant at 60–65 years, and performed differential expression and Gene Ontology (GO) enrichment analyses [39] for these late-changing features in both the accelerated and decelerated aging groups (Methods). For instance, in accelerated aging, enrichment for T-cell activation [40, 41] indicates an adaptive-immune, adhesion-driven inflammatory state that is consistent with chronic antigenic stimulation and tissue infiltration (Fig. 3b). In stark contrast, the decelerated aging group exhibited a blunted and delayed molecular aging phenotype, with a significantly smaller number of features changing in early mid-life while the number of features that change during the transition from middle age to old age increases dramatically (Fig. 3c). Unlike accelerated aging, decelerated aging is marked by enrichment of regulation of small-molecule metabolic process, aligning with preserved metabolic and redox homeostasis that supports genome maintenance and antigen presentation while restraining excessive immune activation[42, 43] (Fig. 3d).
@@ 4.2 Body system-derived features
To systematically analyze the interrelationships among physiological modules, we organized the hundreds of measurement parameters and questionnaire information collected at baseline into the following core system-level categories.
@@
4.3 Transcriptomics
Transcriptomic profiling was conducted on human peripheral blood mononuclear cells (PBMCs). RNA extraction was performed from PBMC samples obtained from participants at the Clinical Testing Center (CTC) following standard operating procedures. Blood was collected in three tubes per participant (Complete Blood Count, serum, and PBMC isolation). PBMCs were processed within 2 hours of collection by centrifugation to isolate the mononuclear layer. Following isolation, PBMCs were aliquoted into three tubes per sample, and one aliquot designated for RNA sequencing was stored at -80°C to preserve RNA integrity. Cell viability and concentration were evaluated using an automated cell counter to ensure high-quality samples for downstream processing. For library preparation, total RNA was extracted using a standard protocol. During reverse transcription, unique molecular identifiers (UMIs), sample barcodes, PCR primers, and ERCC spike-ins were incorporated into each sample. ERCC spike-ins served as external controls for quantification and normalization of gene expression levels. The resulting cDNA underwent initial amplification to increase material yield for library construction. Subsequently, a tagmentation reaction was performed using a cutting enzyme (TDE1, or Nextera transposase prior to May 2023) to fragment the cDNA and introduce sequencing adaptors. Following fragmentation, samples were pooled, and a final round of PCR amplification was carried out to complete the cDNA library for sequencing. RNA sequencing libraries were sequenced on an Illumina NovaSeq platform with an average read depth of approximately 5 million reads per sample. Read 1 (R1) was 20 base pairs in length (containing the UMI and poly-T tail), while Read 2 (R2) was 66 base pairs in length and contained the cDNA fragment and possibly the poly-A tail. PhiX control libraries were spiked into sequencing runs to enhance base diversity. The sequencing generated high-quality 3’ RNA-seq data suitable for downstream bioinformatic analyses.
4.4 Metabolomics
Metabolite profiling was conducted using the proton nuclear magnetic resonance (1H-NMR) platform of Nightingale Health, for which the technical details and relevant epidemiological applications have been previously described [50, 51]. This platform provides simultaneous quantification of 228 absolute-value-based plasma metabolites and ratios, with concentrations estimated from reference data. The profiling primarily extends detailed lipidomic characterization and includes measurements of clinically validated biomarkers, such as routine lipids, lipoprotein subclass profiling with lipid concentrations within 14 subclasses, fatty acid composition, and various low-molecular-weight metabolites including amino acids, ketone bodies, and glycolysis metabolites. For data processing, instances of the value ’TAG’ were replaced with zero, and a corresponding binary indicator variable was added. No further normalization or imputation was applied.
4.5 Serum lipidomics
[…]
4.6 Microbiome
Fecal samples were collected using OMNIgene-Gut stool collection kits (DNA Genotek), and subgingival plaque samples were obtained by a dentist using the same type of kit. Genomic DNA was extracted with the PowerMag Soil DNA Isolation Kit (MoBio), optimized for use on a Tecan automated platform. Shotgun metagenomic sequencing was performed on either the Illumina NextSeq 500 (single-end, 75 bp reads) or NovaSeq 6000 (single-end, 100 bp reads) platforms. Raw sequencing reads were processed with Trimmomatic0.38 for adapter removal and quality filtering using the following parameters: -phred33 ILLUMINACLIP:<<adapter file>>:2:30:10 SLIDINGWINDOW:6:20 CROP:75 MINLEN:65 for 75 bp reads, and CROP:100 MINLEN:90 for 100 bp reads. Reads mapping to the human genome (hg19) were identified with Bowtie2 and excluded from downstream analyses. Species-level composition was determined in two steps. First, non-human reads were aligned with Bowtie2 to a subset of representative genomes from Pasolli et al. [53], limited to species-level genome bins (SGBs) containing at least five genomes. Taxonomic assignments were based on a majority vote among reference genomes within each SGB; when reference genomes were absent, higher-level taxonomic labels were assigned following the criteria in Pasolli et al[53]. In the second step, species relative abundance was calculated using the Unique Relative Abundance (URA) method, which estimates abundance as the mean coverage of the 50% most densely covered genomic regions, considering only uniquely mapped reads. Parameters were as follows: for 75 bp gut samples [min_mapped_to_retain = 1 M reads; num_mapped_to_subsample = 8 M reads], for 75 bp oral samples [0.5 M, 4 M reads], for 100 bp gut [1 M, 5 M reads], and for 100 bp oral [0.5 M, 2.5 M reads]. Shannon $\alpha$-diversity and species richness were computed prior to the data processing steps described above. The proportion of human DNA shedding was estimated as the fraction of human reads removed relative to the total number of reads passing quality control. Microbial functions were calculated using MetaCyc24 pathways by HUMAnN3, Bowtie2, DIAMOND2 and MetaPhlAn4 [54, 55, 56, 57].